A student has test scores of 68, 75 and 79. What must she score on the fourth exam to have an average of at least 80 if the maximum score is 100?
Answer by harpazo(655)


(68+75+79+x)/4 = 80
(222+x)/4 = 80
Multiply both sides by 4.
222+x = 320
x = 320-222
x = 98.
The student must score 98 on the fourth exam.